# Video Understanding with Large Language Models: A Survey

### v5

@@ Abstract
Index Terms:
Video Understanding, Large Language Model, Vision-Language Model, Multimodality Learning
@@ II-B Background for LLMs
Decoding strategies dictate how to harness the next token probability and select the next token $y_{t}$ from the set $S$ of all possible tokens in the vocabulary, which includes special tokens such as $<$SOS$>$, $<$EOS$>$, and $<$PAD$>$. Greedy decoding, the simplest strategy, selects the token with the highest probability, formalized as:
@@ III Vid-LLMs
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=6em, edge+=darkgray, line width=1pt, s sep=20pt, inner xsep=3pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10.2em,font=,, where level=2text width=8.6em,font=,, where level=3text width=8.2em,font=,, where level=4text width=8.2em,font=,, [ Vid-LLMs, ver, my-box, text width=12em [ Video Analyzer $\times$ LLM, analyzer-box, edge=myorange [ LLM as Summarizer, analyzer-box, edge=myorange [ GIT-LLaVA [156], MM-Screenplayer [157], MoReVQA [158],
IG-VLM [159], LangRepo [160], MVU [161],
Video ReCap [162], LLoVi [163], Grounding-Prompter [164],
VIDOSC [165], AntGPT [166], VAST [105],
VLog [167], LaViLa [168] , leaf, text width=27em, edge=myorange, draw=myorange ] ] [ LLM as Manager, analyzer-box, edge=myorange [ DrVideo [169], OmAgent [170], LVNet [171], GPTSee [172],
VideoTree [173], LAVAD [174], TraveLER [175],
RAVA [176], SCHEMA [177], HuggingGPT [178],
TV-TREES [179], VideoAgent [180], VideoAgent [181],
VURF [182], KEPP [183], DoraemonGPT [184], Hawk [185],
LifelongMemory [186], ProViQ [187], AssistGPT [188],
Video ChatCaptioner [189], ChatVideo [189], ViperGPT [190] , leaf, text width=27em, edge=myorange, draw=myorange ] ] ] [ Video Embedder $\times$ LLM, embedder-box, edge=mypink [ LLM as Text
Decoder, embedder-box, edge=mypink [ Artemis [191], EmoLLM [192], FTFV-LLM [193],
Flash-VStream [194], LLAVIDAL [195], LongVA [196],
ShareGPT4Video [197], VIM [198], Video-SALMONN [199],
VideoGPT+ [200], VideoLLaMA 2 [201], MotionLLM [202],
VideoChat2 [203], Shotluck Holmes [204], VideoStreaming [205],
VideoNarrator [206], TOPA [207], AutoAD III [208], GCG [209],
LLaVA-Hound-DPO [210], RED-VILLM [211], Koala [212],
LongVLM [213], MA-LMM [214], MiniGPT4-Video [215],
Pegasus-v1 [216], PLLaVA [217], ST-LLM [218], COSMO [219],
Tarsier [220], X-VARS [221], CAT [222], VideoLLM [223],
InternVideo2 [224], MovieLLM [225], IVAwithLLM [226],
LSTP [227], LVCHAT [228], OSCaR [229], Slot-VLM [230],
AV-LLM [231], Emu2 [232], MMICT [233], VaQuitA [234],
VILA [235], Vista-LLaMA [236], Chat-UniVi [237],
LLaMA-VID [238], Video-LLaVA [239], LLaMA-VQA [240],
MovieChat [241], LLMVA-GEBC [242], Macaw-LLM [243],
VALLEY [244], Video-ChatGPT [99], Video-LLaMA [245],
mPLUG-video [96], ChatBridge [246], Otter [247] , leaf, text width=28em, edge=mypink, draw=mypink ] ] [ LLM as Regressor, embedder-box, edge=mypink [ Holmes-VAD [248], VideoLLM-online [249], VLM4HOI [250],
V2Xum-LLaMA [100], AVicuna [120], Elysium [251],
HawkEye [252], LITA [253], OmniViD [254], SeViLA [255],
GroundingGPT [256], TimeChat [257], VTimeLLM [258] , leaf, text width=27em, edge=mypink, draw=mypink ] ] [ LLM as Hidden
Layer, embedder-box, edge=mypink [ VTG-LLM [259], VITRON [260], VTG-GPT [261],
Momentor [262], VidDetours [263], OneLLM [264],
GPT4Video [265] , leaf, text width=23em, edge=mypink, draw=mypink ] ] ] [ (Analyzer + Embedder)
$\times$ LLM, plus-box, edge=mygreen [ LLM as Manager, plus-box, edge=mygreen [ MM-VID [266] , leaf, text width=7em, draw=mygreen, edge=mygreen ] ] [ LLM as Summarizer, plus-box, edge=mygreen [ SUM-shot [267] , leaf, text width=7em, draw=mygreen, edge=mygreen ] ] [ LLM as Regressor, plus-box, edge=mygreen [ Vriptor [268], Merlin [269], VideoChat [270], Vid2Seq [271] , leaf, text width=26.5em, draw=mygreen, edge=mygreen ] ] [ LLM as Text
Decoder, plus-box, edge=mygreen [ Uni-AD [272], MM-narrator [273], Vamos [274], Auto-AD II [275] , leaf, text width=29em, draw=mygreen, edge=mygreen ] ] [ LLM as Hidden
Layer, plus-box, edge=mygreen [ PG-Video-LLaVA [276] , leaf, text width=10em, draw=mygreen, edge=mygreen ] ] ] ]
@@ V-B3 Multimodal Video Understanding
Multimodal video understanding requires integrating different types of data, such as visual, audio, and text, for a better understanding of videos [120, 367]. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. The field lacks research and datasets, with challenges in ensuring high-quality data annotation.

### v7

@@ II-B Background for LLMs
Decoding strategies dictate how to harness the next token probability and select the next token $y_{t}$ from the set $S$ of all possible tokens in the vocabulary, which includes special tokens such as <<SOS>>, <<EOS>>, and <<PAD>>. Greedy decoding, the simplest strategy, selects the token with the highest probability, formalized as:
@@ III Vid-LLMs
[…]
@@ V-B3 Multimodal Video Understanding
Multimodal video understanding requires integrating different types of data, such as visual, audio, and text, for a better understanding of videos [120, 368, 369]. Aligning these data, especially in terms of spatial and temporal synchronization, is particularly critical. This area lacks relevant research and suffers from a scarcity of datasets. The field lacks research and datasets, with challenges in ensuring high-quality data annotation.